Two angles are complementary if the sum of the measures 90. The measure of the first angle is X and the second angle is 22° less than three times the first. Find measure of the two angles
Answer by Boreal(15235)

x and 90-x are the angles, for they add to 90
second angle 90-x is 3x-22
therefore, 90-x=3x-22
112=4x
x=28 degrees, 90-x=62 degrees, the angles ANSWER
28*3-22=62. checks